OpenStudy (joftheworld):

can someone help me with 8 functional value problems ? ill offer 10 medals and a fan !

OpenStudy (joftheworld):

1st problem

OpenStudy (anonymous):

subsiture where there is x and where there is y. then continue to solve the eqution.

OpenStudy (joftheworld):

i dont get it ..

OpenStudy (anonymous):

so basically... (x,y) . your function says..f(x+2) y-f(0+2) =7

OpenStudy (anonymous):

solve this

OpenStudy (anonymous):

y =*

OpenStudy (amistre64):

without knowing what f(x) defines, then f(0) = 7 tells us nothing about f(0+2)

OpenStudy (amistre64):

the question makes no sense to me without a graph/definition

OpenStudy (amistre64):

for example, the graphs for: g(x) = x+7; h(x) = x^2+7 both contain the point (0,7). f(x+2) can be seen as a shift, unless defined otherwise in your material

OpenStudy (joftheworld):

want me to move on to the next question ?

OpenStudy (amistre64):

if its as badly constructed as this one .... then no :)

OpenStudy (joftheworld):

okay well any idea on how to solve it ?...

OpenStudy (amistre64):

lion suggestion is just as valid as any other that i can think of. the question makes no sense out of context of your material, and very well may make no sense period.

OpenStudy (joftheworld):

exactly why i offered 10 medals lol because i dont get this at all..

OpenStudy (amistre64):

is there a graph for f(x)?

OpenStudy (amistre64):

if so, then f(2) corresponds to the proper solution

OpenStudy (amistre64):

f(0) -> (0,7) f(0+2) -> (0,f(2)) for whatever f(2) is defined for on the graph of f(x)

OpenStudy (joftheworld):

no but you can graph it here i think.. https://www.desmos.com/calculator

OpenStudy (amistre64):

you cant graph it without the function being defined in some manner. like f(x) = x+7, f(0)=7, f(2)=9 or f(x) = x^2+7, f(0) = 7, f(2)=11 or f(x) = sin(x)+7, f(0) = 7, f(2) = 7.91 or f(x) = etc ....

OpenStudy (amistre64):

in all these cases (0,7) is on the graph of some defined f(x), and as we can see, f(0+2) is not the same in any of the cases.

OpenStudy (joftheworld):

hm... so it cant be solved ?

OpenStudy (amistre64):

not in any numerical way that i can see. the corresponding point may just have to be written as (0,f(2)) since there is no way to actually define f(2)

OpenStudy (amistre64):

f(2) may not even exist on f(x)

OpenStudy (amistre64):

f(x) = -14/(x-2), has a point at 0,7 but is undefined at x=2

OpenStudy (amistre64):

go thru your lessons and see if any of it pertains to this question. other than that ...

OpenStudy (joftheworld):

i can just skip this ..